# EmoSphere++: Emotion-Controllable Zero-Shot Text-to-Speech via Emotion-Adaptive Spherical Vector

### v1

@@ III-A Emotion-Adaptive Coordinate Transformation
As Fig. 2 (a) shows, the simplest way to determine the intensity of an emotion $k$ is to assume that the emotion intensity decreases as it moves closer to the neutral emotion center $M$ as follows [3, 12]:
$M=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}e_{i}^{n}.$ (1)
Here, $N_{n}$ is the total number of a neutral coordinate $e_{ni}$. However, as the mean-based method does not consider the distribution of other emotion categories, such as variance, it does not fully capture the information of the target emotion. In our proposal, we simultaneously consider the distribution of the target and other emotion categories and extract an adaptive spherical vector of each emotion category.
We studied a coordinate method to measure the angle-based distance measure, which can model emotional style and intensity on a continuous scale [3, 34]. The key idea of the proposed method is to model a spherical coordinate system for each emotion by simultaneously considering the distribution of the target and other emotion categories, as shown in Fig. 2 (b). We first considered two assumptions: (1) the emotional intensity increases as it moves farther from the center of the emotion-adaptive spherical coordinate system and (2) the angle from the center of the emotion-adaptive spherical coordinate determines the emotional style. Initially, we adopted a specific speech emotion recognition (SER) model [48] $\psi$ to predict the VAD value $e_{i}^{k}$ in emotion class $k$:
@@
where $x_{i}$ denote the $i$-th referece speech in speech dataset $\mathrm{X}$ and $e_{i}^{k}$ consists of values for $(d_{v},d_{a},d_{d})$, where $d_{v}$, $d_{a}$, and $d_{d}$ represent valence, arousal, and dominance, respectively. Each component is expressed in Cartesian coordinates, with values ranging from 0 to 1. To model the spherical coordinate system for each emotion, we obtained the shifted Cartesian coordinates $\widehat{e_{i}^{k}}=(\widehat{d}_{v},\widehat{d}_{a},\widehat{d}_{d})$ by shifting through the representative central coordinates $M_{k}$ from different emotion coordinate set $E_{k}$. The coordinates $M_{k}$ are extracted using emotion-specific centroid extraction, which maximizes the ratio of the distance between specific emotions to the distance from the neutral coordinates as follows:
@@ 20   Append $s_{i}^{k}=(r_{IQR},\vartheta,\varphi)$ to $\mathbb{S}$
Here $e_{i}^{k}$ and $e_{i}^{n}$ denote the $i$-th coordinate of the $k$-th emotion coordinate set $E_{k}$ and the neutral coordinate set $E_{n}$, respectively. Consequently, the centroid coordinates are the ones that maximize the distances from the target emotion category while minimizing the distance within the neutral emotion category. Then, transformation via the representative central coordinates $M_{k}$ to spherical coordinates $s_{i}^{k}=(r,\vartheta,\varphi)$ can be formulated as follows:
@@
III-B Multi-Level Style Encoder
The multi-level style encoder consists of two main modules: the emotion and speaker encoders. The emotion encoder comprises high- and low-level emotion encoders.
As shown in Fig. 3, the emotion encoder includes a fine-tuned SER model for high-level emotion features and an additional EASV extractor for low-level emotion features. The high-level emotion encoder extracts the fixed-size hidden embedding from the SER [^1] https://github.com/ddlBoJack/emotion2vec [50], which contains a multilayer transformer. Meanwhile, the EASV extractor, which is described in Section III-A, generates low-level emotion features. A fully connected layer processes high- and low-level emotion features, combining them into a fixed-size hidden embedding.
The speaker encoder provides speaker-related information to the TTS model. A pretrained speech encoder extracts a speaker embedding [51] from the reference speech for zero-shot emotion transfer. We used the WavLM Base model [52] as the speaker verification model [^2] https://huggingface.co/microsoft/wavlm-base-sv, building the model on the HuBERT [53] framework to focus on modeling speech content and preserving speaker identity. WavLM-based approaches can capture and represent speaker-specific information and emotional nuances in speech [54, 55]. Similar to emotion embeddings, fixed-size hidden embeddings are processed in the fully connected layer and combined to generate multi-level style embedding $e_{sty}$.
@@ III-C Preliminary on Conditional Flow Matching-Based Model
EmoSphere++ adopts a conditional flow matching (CFM)-based decoder that generates flow through the ordinary differential equation. Building on the success of flow matching in the speech synthesis task [23, 25, 24], we utilized a CFM-based decoder that is designed to model a conditional vector field $\mathbf{u}_{t}$. Following [56], we define the flow $\phi$ as the mapping between two density functions:
@@
Here, $\boldsymbol{v}_{t}$ represents a time-dependent vector field that defines the path of the probability flow over time $t\in\left[0,1\right]$. Particularly, it describes a conditional flow process in which the conditional flow $\phi_{t,x1}$ represents simple linear trajectories between the data point $x_{1}$ drawn from the target distribution $q(x)$ and prior distribution $x_{0}\sim N(0,I)$:
@@
where $\mu$ represents the predicted average acoustic features (e.g., Mel-spectrogram) given the text and the chosen durations, using a text encoder and duration predictor. $e_{sty}$ denotes the multi-level style embedding.
@@ III-D Training Objective
Alongside traditional losses for training TTS systems, we introduce objective functions to help in modeling our system to control and transfer emotion style and intensity.
First, text encoder and duration predictor architectures were implemented based on [23]. Duration-model training uses monotonic alignment search to compute the duration loss $L_{dur}$ and the prior loss $L_{enc}$, as described in [57]. The decoder follows the CFM loss $L_{cfm}$ described in Section III-C.
Inspired by [15], we introduced an additional orthogonality loss. In the style encoder, emotion and speaker embeddings contain overlapping information regarding each other. The speaker and emotion embeddings should be 1) discriminative in distinguishing identities and 2) independent of each other, ensuring effective generalization performance for both seen and unseen speakers. To mitigate the impact of speaker and emotion embedding leakage on model performance, we propose a normalized orthogonality loss $L_{ort}$ to enhance the decoupling capability of the model. Unlike the existing loss, this method normalizes all sample pairs, providing effective generalized performance for zero-shot scenarios. In this case, $L_{ort}$ can be expressed as:
@@
Figure 4: Run-time diagram of the proposed EmoSphere++ framework. The emotion-adaptive spherical vector can manually control the low-level emotion of emotion style and intensity. We produce an emotional state as the derivative of primary emotions by assigning the appropriate angle and length to the spherical vector.
@@ III-E Control of Emotional Style and Intensity
Fig. 4 illustrates the proposed emotion-controllable zero-shot TTS framework, which synthesizes emotional speech for both seen and unseen speakers based on reference speech and EASV. The framework comprises three main modules: the text encoder, multi-level style encoder, and CFM decoder.
The multi-level style encoder captures the emotion and speaker information in an embedding from the reference speech. By varying the length and angle in the EASV, we can manipulate the levels of style and intensity at runtime and efficiently synthesize the desired emotional effects. In a zero-shot scenario, the framework can synthesize speech for unseen speakers by inputting the unseen target reference speech into the speaker encoder.
TABLE I: Prosodic variation analysis of emotion-adaptive spherical vector values on ESD datasets. Q1, Q2, and Q3 represent intensity regions divided by 0.33 and 0.66 thresholds. The average prosodic change $p_{c}$ shows the ratio of the average of each style to the overall emotion average. “-” indicates missing data, with neutral emotions fixed at intensity 0. Cell hue represents heat-mapped prosody values for each emotion.
Emotion Style Num. Pitch (mean) Energy (mean) Duration (mean)
Q1 Q2 Q3 All Q1 Q2 Q3 Avg. $p_{c}$ Q1 Q2 Q3 Avg. $p_{c}$ Q1 Q2 Q3 Avg. $p_{c}$
Neutral All (Average) - - - 3,500 - - - 51.68 - - - - 0.41 - - - - 2.60 -
Angry I (+V +A +D) 297 788 137 1,222 55.00 66.02 76.79 64.68 +0.14 0.99 1.43 1.86 1.37 +0.69 2.96 2.85 2.81 2.87 +0.05
IV (+V -A +D) 1 - - 1 47.78 - - 47.78 -0.16 0.67 - - 0.67 -0.18 2.28 - - 2.28 -0.17
V (+V +A -D) 374 1,365 533 2,272 53.87 65.64 78.65 66.92 +0.18 1.05 1.37 1.73 1.40 +0.73 2.79 2.76 2.72 2.76 +0.01
VII (-V -A -D) 3 - - 3 51.96 - - 51.96 -0.08 0.97 - - 0.97 +0.20 2.85 - - 2.85 +0.04
VIII (+V -A -D) 2 - - 2 57.29 - - 57.29 +0.01 0.02 - - 0.02 -0.98 2.65 - - 2.65 -0.03
All (Average) 677 2,153 670 3,500 54.33 65.78 78.26 56.66 - 1.01 1.40 1.76 0.81 - 2.87 2.79 2.74 2.75 -
Happy I (+V +A +D) 338 1,482 472 2,292 64.46 77.01 90.05 77.84 +0.62 0.59 0.83 0.99 0.83 +0.83 2.77 2.67 2.66 2.68 -0.11
II (-V +A +D) 4 5 - 9 35.165 33.56 - 34.24 -0.29 0.23 0.16 - 0.19 -0.58 2.94 3.31 - 3.15 +0.04
III (-V -A +D) 81 18 - 99 31.99 28.20 - 31.37 -0.35 0.21 0.17 - 0.20 -0.56 3.38 3.38 - 3.38 +0.12
IV (+V -A +D) 20 - - 20 52.54 - - 52.54 +0.09 0.39 - - 0.39 -0.13 2.96 - - 2.96 -0.02
V (+V +A -D) 159 666 171 996 62.93 78.66 89.26 78.09 +0.63 0.67 0.90 1.34 0.94 +1.08 2.69 2.58 2.66 2.61 -0.13
VI (-V +A -D) 2 - - 2 31.01 - - 31.01 -0.36 0.08 - - 0.08 -0.83 3.60 - - 3.60 +0.19
VII (-V -A -D) 56 12 - 68 32.28 26.74 - 31.44 -0.35 0.22 0.23 - 0.22 -0.51 3.36 3.07 - 3.31 +0.10
VIII (+V -A -D) 14 - - 14 47.75 - - 47.75 -0.01 0.76 - - 0.76 +0.68 2.41 - - 2.41 -0.19
All (Average) 674 2,183 643 3,500 55.16 76.63 89.84 48.03 - 1.31 1.97 2.51 0.45 - 2.88 2.65 2.66 3.02 -
Sad I (+V +A +D) 13 29 1 43 61.86 76.17 90.43 72.17 +0.13 0.94 1.04 2.00 1.04 +0.27 2.82 2.71 2.92 2.75 -0.05
II (-V +A +D) 1 5 - 6 33.82 54.10 - 51.03 -0.21 0.89 0.65 - 0.70 -0.14 3.89 3.03 - 3.17 +0.1
III (-V -A +D) 184 713 98 995 57.95 53.21 43.51 53.15 -0.17 0.72 0.60 0.38 0.60 -0.27 2.95 3.06 3.25 3.06 +0.06
IV (+V -A +D) 39 11 - 50 72.76 75.94 - 73.42 +0.14 0.88 0.95 - 0.89 +0.09 2.83 2.68 - 2.79 -0.03
V (+V +A -D) 46 113 15 174 68.97 76.37 67.01 73.68 +0.15 1.11 1.03 0.87 1.04 +0.28 2.97 2.78 2.31 2.79 -0.03
VI (-V +A -D) 3 4 - 7 55.81 65.45 - 62.01 -0.03 0.67 0.67 - 0.67 -0.18 2.60 3.02 - 2.84 -0.02
VII (-V -A -D) 228 1,336 535 2,099 62.74 53.83 38.81 51.10 -0.20 0.70 0.67 0.52 0.64 -0.22 2.91 2.93 3.15 2.98 +0.03
VIII (+V -A -D) 54 70 2 126 74.43 78.60 82.04 76.84 +0.20 0.99 0.93 0.88 0.96 +0.17 2.67 2.71 2.98 2.70 -0.07
All (Average) 568 2,281 651 3,500 63.36 55.87 40.28 64.18 - 0.79 0.68 0.51 0.82 - 2.89 2.95 3.15 2.89 -
Surprise I (+V +A +D) 339 1,083 279 1,693 61.13 83.05 93.42 80.48 +0.35 0.98 1.33 1.88 1.35 +0.15 2.87 2.75 2.76 2.78 -0.01
IV (+V -A +D) 3 - - 3 46.2 - - 46.2 -0.23 0.75 - - 0.75 -0.36 3.5 - - 3.5 +0.25
V (+V +A -D) 322 1,079 385 1,786 60.68 82.14 103.48 82.89 +0.39 1.01 1.33 2.09 1.43 +0.22 2.77 2.7 2.56 2.69 -0.04
VII (-V -A -D) 1 - - 1 48.73 - - 48.73 -0.19 1.35 - - 1.35 +0.16 2.49 - - 2.49 -0.11
VIII (+V -A -D) 9 - - 9 50.14 - - 50.14 -0.16 1.26 - - 1.26 +0.08 2.53 - - 2.53 -0.1
All (Average) 674 2,162 664 3,500 60.59 82.60 99.16 59.77 - 0.99 1.33 2.00 1.17 - 2.82 2.73 2.65 2.80 -
@@ IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^3] https://github.com/HLTSingapore/Emotional-Speech-Data[58], which contains 350 parallel utterances spoken by 10 English speakers in five emotional states (neutral, happy, angry, sad, and surprise). Following the prescribed data partitioning criteria, we extracted one sample for each emotion from every speaker, resulting in 17,500 samples. The validation set comprised 20 samples for each emotion per speaker, totaling 1,000 samples, whereas the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. The zero-shot scenario used two unseen speakers, one English-speaking male (“0013”) and one English-speaking female (“0019”), by excluding them from the training process.
@@
Moreover, we conducted experiments on the interactive emotional dyadic motion capture database (IEMOCAP) [59] to verify whether the models can reflect the styles using predicted VAD values. The IEMOCAP contains 10 speakers with nine emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral) and dimensional labels such as VAD. The validation set included 10 samples for each speaker, whereas the test set comprised 15 samples.

### v2

@@ III-A Emotion-Adaptive Coordinate Transformation
Several studies [12, 7, 10, 3] assume that emotional intensity decreases when approaching a neutral state and use this as the basis for modeling emotion intensity. As shown in Fig. 2 (a), previous studies [3] defined the center coordinates based on this assumption, where the intensity of emotion decreases as it approaches the neutral emotion center $M$, formulated as follows:
$M=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}e_{i}^{n},$ (1)
where $N_{n}$ is the total number of neutral coordinates $e_{i}^{n}$. However, since the mean-based method does not account for the distribution of other emotions, such as variance, it fails to fully capture the relationship between the neutral emotion and the target emotion. Specifically, we define other emotions as all emotions except the neutral emotion and the target emotion as a specific emotion selected from these other emotions. To address this, our approach simultaneously considers the distributions of the neutral emotion and the corresponding target emotion, extracting an adaptive spherical vector for each target emotion.
Our method models a spherical coordinate system for each target emotion by considering the distribution of the neutral emotion and its corresponding target emotion, as shown in Fig. 2 (b). Our approach is based on two fundamental assumptions: (1) the emotional intensity increases as it moves farther from the center of the emotion-adaptive spherical coordinate system and (2) the angle from the center of the emotion-adaptive spherical coordinate determines the emotional style. Initially, we adopted a specific emotional attribute prediction model [53] $\psi$ to predict the VAD value $e_{i}^{k}$ in emotion class $k$:
@@
where $x_{i}$ denote the $i$-th referece speech in speech dataset $\mathrm{X}$ and $e_{i}^{k}$ consists of values for $(d_{v},d_{a},d_{d})$, where $d_{v}$, $d_{a}$, and $d_{d}$ represent valence, arousal, and dominance, respectively. Each component is expressed in Cartesian coordinates, with values ranging from 0 to 1. To model the spherical coordinate system for each emotion, we obtained the shifted Cartesian coordinates $\widehat{e_{i}^{k}}=(\widehat{d}_{v},\widehat{d}_{a},\widehat{d}_{d})$ by shifting through the representative central coordinates $M_{k}$ from different target emotion coordinate set $E_{k}$. The coordinates $M_{k}$ are extracted using emotion-specific centroid extraction, which maximizes the ratio of the distance between specific target emotions to the distance from the neutral coordinates as follows:
@@ 20   Append $s_{i}^{k}=(r_{IQR},\vartheta,\varphi)$ to $\mathbb{S}$
Here $e_{i}^{k}$ and $e_{i}^{n}$ denote the $i$-th coordinate of the $k$-th target emotion coordinate set $E_{k}$ and the neutral coordinate set $E_{n}$, respectively. Consequently, the centroid coordinates are the ones that maximize the distances from the target emotion category while minimizing the distance within the neutral emotion category. Then, transformation via the representative central coordinates $M_{k}$ to spherical coordinates $s_{i}^{k}=(r,\vartheta,\varphi)$ can be formulated as follows:
@@
III-B Joint Attribute Style Encoder
The voice typically contains highly dynamic style attributes (e.g. speaker identities, prosody and emotions), making the TTS model difficult to model and transfer in a zero-shot scenario. As shown in Fig. 3, we propose a joint attribute style encoder for both broad and fine-grained stylization.
The emotion encoder includes a fine-tuned categorical emotion recognition model for global emotion features and an additional EASV extractor for dimensional-driven emotion features. The global emotion encoder extracts the fixed-size hidden embedding from the categorical emotion recognition [^1] https://github.com/ddlBoJack/emotion2vec [55], which utilizes a multilayer transformer to capture comprehensive emotional representations. Meanwhile, the EASV extractor, described in Section III-A, generates dimensional-driven emotion features, which explicitly encode fine-grained variations in emotional expressions (e.g., intensity and style). Finally, a fully connected layer processes global and dimensional-driven emotion features, combining them into a fixed-size hidden embedding.
The speaker encoder provides speaker-related information to the TTS model. A pretrained speech encoder extracts a speaker embedding [56] from the reference speech for zero-shot emotion transfer. We used the WavLM Base model [57] as the speaker verification model [^2] https://huggingface.co/microsoft/wavlm-base-sv, building the model on the HuBERT [58] framework to focus on modeling speech content and preserving speaker identity. WavLM-based approaches can capture and represent speaker-specific information and emotional nuances in speech [59, 60]. Similar to emotion embeddings, fixed-size hidden embeddings are processed in the fully connected layer and combined to generate joint attribute style embedding $e_{sty}$.
@@ III-C Preliminary on Conditional Flow Matching-Based Model
EmoSphere++ adopts a CFM-based decoder that generates flow through the ordinary differential equation. Building on the success of flow matching in the speech synthesis task [26, 28, 27], we utilized a CFM-based decoder that is designed to model a conditional vector field $\mathbf{u}_{t}$. Following [61], we define the flow $\phi$ as the mapping between two density functions:
@@
Here, $\boldsymbol{v}_{t}$ represents a time-dependent vector field that defines the path of the probability flow over time $t\in\left[0,1\right]$. Specifically, it describes a conditional flow process in which the conditional flow $\phi_{t,x1}$ represents simple linear trajectories between the data point $x_{1}$ drawn from the target distribution $q(x)$ and prior distribution $x_{0}\sim N(0,I)$:
@@
where $\mu$ represents the predicted average acoustic features (e.g., Mel-spectrogram) given the text and the chosen durations, using a text encoder and duration predictor. $e_{sty}$ denotes the joint attribute style embedding.
@@ III-D Training Objective
Alongside traditional losses for training TTS systems, we introduce a disentanglement method to help in modeling our system to control and transfer emotion style and intensity.
First, text encoder and duration predictor architectures were implemented following [26]. Duration-model training uses monotonic alignment search to compute the duration loss $L_{dur}$ and the prior loss $L_{enc}$, as described in [62]. The decoder follows the CFM loss $L_{cfm}$ described in Section III-C.
Inspired by [15], we introduced an additional orthogonality loss of the disentanglement method. In the style encoder, emotion and speaker embeddings contain overlapping information regarding each other. The speaker and emotion embeddings should be 1) discriminative in distinguishing identities and 2) independent of each other, ensuring effective generalization performance for both seen and unseen speakers. To mitigate the impact of speaker and emotion embedding leakage on model performance, we propose a normalized orthogonality loss $L_{ort}$ to enhance the decoupling capability of the model. Unlike the existing loss [15] applied only to embeddings from the same audio, this method normalizes all sample pairs to enhance generalization in zero-shot scenarios. In this case, $L_{ort}$ can be expressed as:
@@
Figure 4: Run-time diagram of the proposed EmoSphere++ framework. We can manually control the emotion style and intensity via the dimensional-driven emotion of emotion style and intensity. We produce an emotional state as the derivative of primary emotions by assigning the appropriate angle and length to the spherical vector.
@@ III-E Control of Emotional Style and Intensity
Fig. 4 illustrates the proposed emotion-controllable zero-shot TTS framework, which synthesizes emotional speech for both seen and unseen speakers based on reference speech and EASV. The framework comprises three main modules: the text encoder, joint attribute style encoder, and CFM decoder.
The joint attribute style encoder captures the emotion and speaker information in an embedding from the reference speech. By varying the length and angle in the EASV, we can manipulate the levels of style and intensity at runtime and efficiently synthesize the desired emotional effects. In a zero-shot scenario, the framework can synthesize speech for unseen speakers by inputting the unseen target reference speech into the speaker encoder.
@@ IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^3] https://github.com/HLTSingapore/Emotional-Speech-Data[63], which contains 350 parallel utterances spoken by ten English speakers in five emotional states (neutral, happy, angry, sad, and surprise). Following the prescribed data partitioning criteria, we extracted one sample for each emotion from every speaker, resulting in 17,500 samples. The validation set comprised 20 samples for each emotion per speaker, totaling 1,000 samples, whereas the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. The zero-shot scenario used two unseen speakers, one English-speaking male (“0013”) and one English-speaking female (“0019”), by excluding them from the training process.
@@
Moreover, we utilized the MSP-Podcast corpus dataset [64] and the Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [65], along with the ESD dataset, to analyze the prosodic variation of EASV and to verify whether the models can reflect the styles using predicted VAD values. The MSP-Podcast corpus dataset comprises approximately 237 hours of speech data annotated with both categorical emotion labels and dimensional VAD values. The training set includes eight categorical emotion classes (happiness, sadness, fear, surprise, contempt, disgust, and neutral) collected from 454 speakers. For dimensional emotion labels, raters evaluated VAD using a seven-point Likert scale. The IEMOCAP contains ten speakers with nine emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral) and dimensional labels such as VAD. For analyzing the prosodic variation of EASV, we used the MSP-Podcast and IEMOCAP datasets while ensuring consistency with ESD by selecting only five categorical emotion labels that match those in the ESD dataset. Additionally, we used the IEMOCAP dataset to verify whether the models can reflect styles using predicted VAD values. The validation set included ten samples per speaker, while the test set comprised 15 samples per speaker.
TABLE I: Prosodic variation analysis of emotion-adaptive spherical vector values across the ESD, IEMOCAP, and MSP-Podcast emotional speech datasets. R1, R2, and R3 represent intensity regions divided by 0.33 and 0.66 thresholds. Green and red colors indicate the highest and lowest prosodic values within each intensity level, respectively. The variation ranges Rc shows the difference between the highest and lowest prosodic values. “-” indicates data not applicable , with neutral emotions fixed at intensity 0.
Emotion Style Num. Pitch (mean) Energy (mean) Duration (mean)
R1 R2 R3 All R1 R2 R3 Rc AVG R1 R2 R3 Rc AVG R1 R2 R3 Rc AVG
Neutral All (Average) - - - 41,845 - - - - 48.7 - - - - 3.2 - - - - 3.7
Angry I (+V +A +D) 611 2,558 480 3,649 66.5 72.9 79.0 12.5 72.5 6.0 7.7 10.9 4.9 7.8 4.1 4.0 3.7 0.4 4.0
III (-V -A +D) 529 2,675 1,123 4,327 59.7 56.1 41.7 18.0 54.4 4.8 2.8 0.8 4.0 2.7 3.9 3.0 1.8 2.1 2.8
IV (+V -A +D) 169 144 6 319 71.1 69.7 64.9 6.2 70.4 8.9 6.4 5.3 3.6 5.9 3.8 3.6 2.6 1.2 3.7
V (+V +A -D) 729 3,702 923 5,354 66.1 72.8 82.6 16.5 73.5 5.7 7.2 15.1 9.4 8.3 4.2 3.8 3.8 0.4 3.8
VI (-V +A -D) 126 107 7 240 55.8 50.9 38.0 17.8 53.4 5.2 4.6 3.4 1.8 4.9 3.9 3.3 2.4 1.5 3.6
VII (-V -A -D) 530 1,763 648 2,941 58.1 54.0 37.3 20.8 53.1 4.8 2.9 0.6 4.2 3.0 3.9 2.5 1.6 2.3 2.6
VIII (+V -A -D) 138 51 - 189 69.9 72.4 - 2.5 70.5 5.5 5.1 - 0.4 5.4 3.9 3.3 - 0.3 3.7
All (Average) 2,832 11,000 3,187 17,019 63.9 64.1 57.3 - 64.0 5.8 5.2 6.0 - 5.4 4.0 3.4 2.7 - 3.5
Happy I (+V +A +D) 825 5,326 1,624 7,775 58.9 64.1 72.2 13.3 65.1 5.2 6.2 8.7 3.5 6.6 4.4 4.1 3.9 0.5 4.1
II (-V +A +D) 233 382 49 664 54.0 49.7 46.3 7.7 50.9 3.5 3.9 3.6 0.4 3.8 4.2 4.3 3.9 0.4 4.2
III (-V -A +D) 1,104 3,715 330 5,149 52.5 46.7 36.4 16.1 47.5 4.0 3.1 1.5 2.5 3.2 4.6 4.3 3.7 0.9 4.4
IV (+V -A +D) 328 360 6 694 61.0 59.0 43.7 17.3 59.9 4.9 4.2 2.8 2.1 4.5 4.3 3.7 4.0 0.3 4.0
V (+V +A -D) 627 2,757 1,568 4,952 61.4 68.3 74.4 13.0 68.7 5.2 5.6 6.6 1.4 5.8 4.1 3.3 2.4 1.7 3.2
VI (-V +A -D) 225 376 149 750 54.0 59.0 56.7 5.0 56.6 3.4 2.5 1.8 1.6 2.8 3.5 2.2 1.7 1.8 2.5
VII (-V -A -D) 866 4,306 1,527 6,699 54.7 50.6 42.4 12.3 49.9 4.1 3.1 1.7 2.4 3.0 4.4 3.7 2.9 1.5 3.6
VIII (+V -A -D) 231 168 18 417 63.7 68.5 69.5 5.8 65.6 5.2 4.6 1.9 3.3 4.9 4.1 3.5 1.7 2.4 3.7
All (Average) 4,439 17,390 5,271 27,100 57.5 58.2 55.2 - 58.0 4.4 4.2 3.6 - 4.3 4.2 3.6 3.0 - 3.7
Sad I (+V +A +D) 412 2,142 810 3,364 47.7 52.7 61.9 14.2 54.6 2.0 2.8 3.9 1.9 3.0 3.5 3.9 4.3 0.8 3.9
II (-V +A +D) 146 146 7 299 44.0 44.4 39.7 4.7 44.1 1.3 1.3 1.1 0.2 1.3 3.0 2.9 2.1 0.9 2.9
III (-V -A +D) 464 1,882 285 2,631 48.2 41.1 30.8 17.4 41.3 1.6 1.2 1.0 0.6 1.2 3.0 2.9 3.1 0.1 2.9
IV (+V -A +D) 111 95 3 209 51.3 51.4 24.8 26.6 50.8 1.7 2.1 1.6 0.5 1.9 3.0 3.5 4.6 1.6 3.2
V (+V +A -D) 419 1,797 589 2,805 50.9 56.1 65.2 14.3 57.5 2.4 3.0 4.4 2.0 3.2 3.1 3.3 3.6 0.5 3.4
VI (-V +A -D) 120 132 7 259 42.6 45.2 36.0 9.2 43.8 1.3 1.8 0.3 1.5 1.5 3.0 2.7 1.8 1.2 2.8
VII (-V -A -D) 411 2,138 891 3,440 48.8 43.5 33.0 15.8 41.8 1.6 1.1 0.7 0.9 1.0 2.8 2.6 2.8 0.2 2.7
VIII (+V -A -D) 115 84 1 200 53.2 63.9 97.6 44.4 57.5 1.3 1.5 0.3 1.0 1.3 2.7 2.4 1.1 1.6 2.5
All (Average) 2,198 8,416 2,593 13,207 48.3 49.8 48.6 - 48.9 1.7 1.8 1.7 - 1.8 3.0 3.0 2.9 - 3.0
Surprise I (+V +A +D) 258 946 423 1,627 71.2 72.3 71.3 1.1 71.9 3.4 4.3 6.6 3.2 4.8 2.4 3.0 3.4 1.0 3.0
III (-V -A +D) 233 1,231 283 1,747 61.7 50.8 41.9 19.8 50.5 3.7 3.3 2.7 1.0 3.2 3.0 3.8 4.1 1.1 3.7
IV (+V -A +D) 61 89 16 166 67.8 64.4 72.2 7.8 66.4 4.1 5.0 2.4 2.6 4.4 3.0 3.5 3.6 0.6 3.3
V (+V +A -D) 256 1,298 437 1,991 71.4 79.9 84.5 13.1 79.7 3.0 3.6 5.1 2.1 3.8 2.4 2.3 2.2 0.2 2.3
VII (-V -A -D) 252 1,052 229 1,533 59.2 54.8 45.1 14.1 54.4 3.6 2.8 1.6 2.0 2.7 2.8 2.6 2.4 0.4 2.6
VIII (+V -A -D) 63 71 5 139 76.0 89.9 84.0 13.9 82.7 4.1 2.3 3.6 1.8 3.2 2.4 1.8 1.6 0.8 2.1
All (Average) 1,123 4,687 1,393 7,203 67.9 68.7 66.5 - 67.6 3.7 3.6 3.7 - 3.7 2.7 2.8 2.9 - 2.8
@@
IV-B Implementation Details
For the acoustic model, we followed the Matcha-TTS [26] configuration, utilizing the text encoder and duration predictor in the encoder, along with a 1D U-Net based CFM decoder. The CFM decoder consists of two downsampling blocks, followed by two midblock and two upsampling blocks, each containing a Transformer layer with a hidden dimensionality of 256, an attention module with dimensionality of 64, and “snakebeta” activations [67]. Following [68], the text encoder was modified by incorporating relative position representations and adding a residual connection to the encoder pre-net. Based on [69], the duration predictor consists of two convolutional layers with rectified linear unit activation, followed by layer normalization, dropout, and a projection layer. Regarding the emotional attribute prediction444huggingface.co/wav2vec2-large-robust-12-ft-emotion-msp-dim, we adopt a system proposed in [53], which predicts VAD using wav2vec 2.0 [70] and a linear predictor. In the joint attribute style encoder, the emotion and speaker global encoders utilize emotion2vec model  [^5] https://github.com/ddlBoJack/emotion2vec [55] and the WavLM-based speaker verification model [^6] https://huggingface.co/microsoft/wavlm-base-sv, respectively. During training, both global encoders are frozen and extract hidden embeddings, which are then passed through a two-layer fully connected network. We trained the generator using random segments of 32 frames from the Mel-spectrogram, with a batch size of 32 and a total of 11M training steps. The AdamW optimizer was used, with a $1\times 10^{-4}$ learning rate. In the inference stage, the guidance level $\gamma$ was set to 100. We trained the vocoder using the official BigVGAN [^7] https://github.com/NVIDIA/BigVGAN[67] implementation, incorporating LibriTTS [71], voice cloning toolkit [^8] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All comparison models were trained using a single NVIDIA RTX A6000 GPU.
IV-C Evaluation
IV-C1 Subjective Metrics
We conducted a MOS evaluation for naturalness (nMOS), speaker similarity (sMOS), and emotion similarity (eMOS) using a nine-point scale ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval of 95%. 20 subjects evaluated the samples by assessing the full set of extracted samples, where two samples were randomly selected for each emotion and speaker combination from the entire test set of 100 samples (2 × 5 (# of emotions) × 10 (8 seen speakers and 2 unseen speakers)), ensuring consistency across models.
We also conducted a preference test to evaluate the modeling of emotion expressiveness. To demonstrate the success of our modeling, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong) from pairs of speech that share the same emotion and style. Evaluators were presented with two different sentences with the same emotion and style, each with varying intensities, and tasked with selecting the one exhibiting a stronger emotion. We uniformly referred to scores 0.1 as weak, 0.5 as medium, and 0.9 as strong. 20 subjects evaluated the samples by assessing the full set of extracted paired samples, where two pairs (four samples) were randomly selected for each emotion and speaker combination from a total of 200 samples in the test set (2 × 2 (pairs) × 5 (# of emotions) × 10 (8 seen speakers and 2 unseen speakers)), ensuring consistency across models.
IV-C2 Objective Metrics
[…]
IV-D Comparison Models
  • GT and BigVGAN [67]: The ground truth (GT) audio and waveforms are generated from the ground truth Mel-spectrogram using vocoder.
  • Mellotron [46]: This auto-regressive multi-speaker TTS model allows for direct style control by conditioning on rhythm and pitch utilizes conditioning GSTs.
  • Mixedemotion [10]: A relative attribute ranking-based model that pre-computes intensity values for mixed emotion synthesis and allows manual control at run time.
  • YourTTS [48]: A zero-shot multi-speaker TTS with the pre-trained speaker encoder. Unlike other comparison models, it does not use a pre-trained vocoder through end-to-end training.
  • GenerSpeech [49]: A high-fidelity zero-shot style transfer method for out-of-domain TTS. We used the emotion2vec model instead of the fine-tuned wav2vec 2.0 model to capture global style for a fair comparison.
  • iEmoTTS [47]: A non-autoregressive TTS model for a cross-speaker emotion transfer system developed based on timbre-prosody disentanglement.
  • EmoSphere++: Our proposed CFM-based emotion-controllable zero-shot TTS with EASV.